OpenStudy (anonymous):

A number added to itself twelve times is five more than twice the number. What is the number?

OpenStudy (cwrw238):

12x = 2x + 5 solve for x

OpenStudy (anonymous):

Can you write the eqn?

OpenStudy (anonymous):

omg howd u figure it out

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

i got 0.5

OpenStudy (cwrw238):

let number be x x + x + x + .. 12 times = 12*x or 12x twice the number = 2x add 5

OpenStudy (cwrw238):

yes that right

OpenStudy (anonymous):

thnk u

OpenStudy (cwrw238):

yw